OpenStudy (anonymous):

Write a two-column proof. Given: 7y = 8x – 14; y = 6 Prove: x = 7

OpenStudy (mathstudent55):

1. 7y = 8x - 14 1. Given 2. y = 6 2. Given 3. 7(6) = 8x - 14 or 3. Substitution 42 = 8x - 14 4. 56 = 8x 4. Addition prop of equality 5. 7 = x 5. Division prop of equality 6. x = 7 6. Symmetric prop of equality

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

wlcm